A popular brand of cola contains 7.80 g of carbon dioxide dissolved in 1.00 L of soft drink. If the evaporating carbon dioxide is trapped in a cylinder at 1.00 atm and 18.0°C, what volume does the gas occupy?
Answer #1

Hello and thank you for your question.

Let's assume the carbon dioxide an ideal gas, and that the whole 7.80 g have evaporated into the cylinder.

Let's recall the ideal gas formula:


PV=nRT

We clear volume:


V=\frac{nRT}{P}

n is the number of mol of the substance we have.

We can determine it as:


n=\frac{mass}{molar mass}

We can find the molar mass of the carbon dioxide on charts.

n=\frac{7.8g}{44.01g/mol}\approx 0.177mol

Now for the constant R is already given as:


R=0.082L.atm/mol.K

We just have to make one small change to apply the formula, convert Celsius into Kelvin:


18^{\circ}C+273.15=291.15K

We proceed to apply the formula:


V=\frac{(0.177mol)(0.082L.atm/mol.K)(291.15K)}{1.00 atm}\approx 4.23L

Good luck!